Soybean meal is 16% protein and Cornmeal is 8%. How many pounds of each should be mixed together in order to get 320 pound mixture that is 15% protein?
Answer by mananth(16949)

protein percent ----------------quantity
Soyabean 16 ----------------x pounds
Cornmeal 8 ---------------- 320 - x pounds
Mixture 15.00% ---------------- 320
16x+8(320-x)=15*320
16x+2560-8x=4800
16x-8x=4800-2560
8x=2240
/8
x=280 pounds Soyabean
40 pounds Cornmeal